The basal energy requirement B is the number of calories that a person needs to maintain the life process. For a 28 year old female with a height of 160 centimenters and a weight of 45 kilograms (kg), B is 1300 calories. If her weight increases to 50 kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Find the equation and find the basal energy requirement if her weight is 53.2 kg.
Answer by stanbon(48568)

For a 28 year old female with a height of 160 centimenters and a weight of 45 kilograms (kg), B is 1300 calories. If her weight increases to 50 kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Find the equation and find the basal energy requirement if her weight is 53.2 kg.
--------------
You have two points:(wt,basal energy)
They are (45,1300), (50,1365)
Find the slope: m = 65/5 = 13
Form of the line is y = mx +b
where m=13 and y=1300 when x=45
1300 = 13*45 + b
b = 715
-----------
EQUATION:
B = 13w + 715
-----------
If her weight is 53.2 kg the basal energy is
B = 13*54.2 + 715
B = 1419.6
================
Cheers,
Stan H.